# Connection between galaxy morphology and dark-matter halo structure II: predicting disk structure from dark-matter halo properties

### v1

@@ 3.2.1 Disk size
Feature importance: The importance rankings differ between the hydro and DMO cases. In the hydro case, the Einasto shape index $\alpha$ is the dominant factor for disk size, while it does not even appear in the top 10 in the DMO case. Some features are consistently influential for both hydro and DMO halos, including concentration $c$, spin $\lambda$, virial mass $M_{\rm vir}$, redshift $z$, and inner-halo spin $\lambda_{\rm inner}$. The most important halo properties align with those used in previous disk-size models (e.g. Mo98; Jiang19). Inner-halo parameters frequently emerge as important features when using the hydro halos, whereas global halo parameters tend to be more important when using the DMO halos. This difference arises mainly because the disk interacts via gravity with the surrounding DM that constitutes the inner halo, altering its structure, as will be further illustrated in Section 4.5. Halo shape parameters are always among the top 10 features. As discussed in Jiang25 and Liang25, this is likely because rounder shapes are indicative of more relaxed halos in which extended disks can more easily develop. Environment and assembly-history parameters (e.g., $n_{1}$, PC1, PC2, $\dot{M}_{\rm norm}$) have limited impact.
@@
3.2.2 Disk scale height
@@
3.2.3 Disk mass fraction
@@
Figure 6: Comparison of models relating disk size to halo properties. Disk half-mass radii $R_{\rm 1/2,Disk}$ measured from the TNG50 hydro simulation are compared with predictions from several models.The first panel shows our SR-based model, written for simplicity as $R_{\rm 1/2,Disk}=f(X)R_{\rm vir}$, where $f(X)$ denotes the SR formula as a function of halo parameters $X$. The cyan and red curves correspond to the SR relations calibrated on hydro halos and DMO halos, respectively, with the DMO version given by eq. (16) and the hydro version by $f(X)=\lambda_{\rm inner}+(\lambda+c^{-0.125})(\log\alpha+3.010)-3.210$. The second panel shows the relation of Kravtsov13, $R_{1/2}=0.02R_{\rm vir}$.The third panel shows a simplified version of the spin-based model of Mo98, $R_{1/2}=0.5\lambda R_{\rm vir}$.The fourth panel shows the concentration-based empirical model of Jiang19, $R_{1/2}=0.02(1+z)^{-0.2}(c/10)^{-0.7}R_{\rm vir}$. Median relations using hydro (DMO) halos are shown as solid cyan (dashed red) lines, with shaded regions indicating the 16–84th percentiles. Spearman correlation coefficients are listed in each panel, and the diagonal line marks the one-to-one relation. This analysis is limited to the galaxies with significant disks ($f_{\rm Disk}>0.3$).  Our SR models provide the highest correlation and lowest scatter, outperforming all previous prescriptions.
@@ 4 Discussion
4.1 New size predictors based on symbolic regression
Simple analytic prescriptions have long been used to link galaxy size to halo properties. Kravtsov13 derived a linear relation $R_{1/2}=0.02\,R_{\rm vir}$ from abundance matching, applicable to both disks and spheroids. Mo98 modeled thin exponential disks that inherit the halo’s specific angular momentum, yielding a size–spin scaling. Following Somerville18, we adopt a much simplified form of the original Mo98 model, $R_{1/2}=0.5\,\lambda\,R_{\rm vir}$. In contrast, Jiang19 found marginal correlation between halo spin and galaxy size in cosmological zoom-in simulations and proposed a concentration-based scaling, $R_{1/2}=0.02(1+z)^{-0.2}(c/10)^{-0.7}R_{\rm vir}$.
@@ 4.1 New size predictors based on symbolic regression
4.2 Understanding the relation between disk size and Einasto shape index
@@
In Section 3, the SHAP analysis shows that the density–profile shape index, $\alpha$, is a key predictor of disk compactness, $R_{1/2,\rm Disk}/R_{\rm vir}$ – but only when using halo properties measured from the hydrodynamic simulation. In the DMO case, $\alpha$ carries far less predictive power. This contrast suggests that the apparent link between disk size and $\alpha$ arises from baryonic effects, rather than from the primordial structure of DMO halos. This is further explored in Fig. 7, which shows a clear positive correlation between $R_{1/2,\rm Disk}/R_{\rm vir}$ and $\alpha$ in the hydro halos, but almost no correlation or weakly negative correlation in the DMO counterparts.
@@
which quantifies the halo’s structural response to baryonic processes. Here, $\alpha_{\rm Hydro}$ is the Einasto index measured in the hydro halo, and $\alpha_{\rm DMO}$ is that of the matched DMO halo. By construction, $\tilde{\alpha}<0$ corresponds to halo contraction (Blumenthal86; Gnedin04), whereas $\tilde{\alpha}>0$ indicates expansion. We also examine the stellar-to-halo mass ratio, $M_{\star}/M_{\rm vir}$, as a proxy for the strength of baryonic impact (DC14; Freundlich20). As shown in the right panels of Fig. 7, most halos have $\tilde{\alpha}<0$, demonstrating that contraction is the dominant response in TNG50. Moreover, the mean value of $\tilde{\alpha}$ decreases with increasing $M_{\star}/M_{\rm vir}$, indicating that halos in more massive galaxies tend to contract more strongly. At fixed stellar-to-halo mass ratio – for example, around $M_{\star}/M_{\rm vir}\simeq 0.03$, where systems are comparably disk-dominated – we find a clear trend that disk compactness governs the degree of halo contraction. The most extended disks (the yellow points in Fig. 7) can even drive halo expansion rather than contraction.
4.3 Redshift trend of disk compactness
@@
To understand these redshift trends, we first focus on the low-mass regime and ask what drives the strong evolution with redshift. A simple and intuitive picture emerges when one approximates the disk size as the radius at which gas first crosses the star-formation density threshold, and considers how the virial radius evolves at fixed halo mass. In the TNG model, star formation begins once the hydrogen number density exceeds $0.13\mathrm{cm}^{-3}$. Using a toy model with an NFW halo and a self-similar gas distribution containing the cosmic baryon fraction, we find that this star-forming radius remains roughly constant across redshift, primarily because typical halo concentrations decrease toward high $z$. By contrast, the virial radius, $R_{\rm vir}\propto(M_{\rm vir}/\rho_{\rm crit})^{1/3}$, becomes much smaller at early times due to the rapid rise of the critical density. As a result, even if the physical disk size evolves only weakly, the ratio $R_{1/2,\rm Disk}/R_{\rm vir}$ increases steeply toward high redshift. This picture is most applicable to low-mass halos, which predominantly grow via cold-mode accretion (Dekel09), where cooling is efficient and disk size closely tracks the extent of star-forming gas.
For more massive systems ($M_{\rm vir}\gtrsim 2\times 10^{11}M_{\odot}$), the processes that shape disk size differ markedly with redshift. At high $z$, these galaxies experience rapid mass assembly and frequent gas-rich mergers, which trigger centrally concentrated starbursts and produce compact stellar components (Zolotov15; Lapiner23). Although often described as star-forming nuggets with spheroidal morphologies, these systems can still retain substantial angular momentum inherited from merger orbits or inflowing filaments. Because our morphological decomposition identifies disks via the angular-momentum distribution of stellar particles, rapidly rotating nuggets are naturally classified as disks. This is not specific to our method – most kinematics-based decomposition schemes behave similarly (e.g., Du19; Zana22) – reflecting the broader fact that high-$z$ disks and fast rotators are not cleanly separable. Consequently, the rotating stellar components in massive high-redshift halos tend to be compact.
At low redshift ($z\lesssim 1$), by contrast, galaxy growth is more quiescent and proceeds inside-out (Genel18). Halos of a given mass are also more dynamically settled, providing favorable conditions for the formation of extended disks (Jiang25). This naturally explains the monotonic increase of disk size with halo mass at low redshift.
4.4 Drivers of disk thickness: mass, size, and halo growth history
As shown in Fig. 4 and discussed in Section 3.2.2, halo properties exhibit unexpectedly strong predictive power for disk thickness, stronger, in fact, than for disk compactness. Here we explore clues to the physical origin of this behavior.
@@ 4.4 Drivers of disk thickness: mass, size, and halo growth history
The disk flattening ratio, $Z_{\rm 1/2,Disk}/R_{\rm 1/2,Disk}$, reaches a minimum of $\sim 0.1$ near $M_{\rm vir}\sim 10^{12}M_{\odot}$ and rises at both lower and higher masses. At any fixed mass, disks are systematically thicker at earlier cosmic times. This reflects enhanced turbulence and elevated vertical velocity dispersion in high-redshift galaxies, consistent with vertical Jeans equilibrium, where $Z_{1/2}\propto\sigma_{z}^{2}/(\pi G\Sigma)$.
@@
A second key finding is the tight correlation between disk height and disk size. As shown in the middle and right panels of Fig. 9, disk height increases almost linearly with disk size at fixed halo mass. This implies that many halo properties influence disk thickness primarily through their regulation of disk size—so the factors that set disk extent also tend to set disk thickness.
@@
We then examine in Fig. 10 the correlations of thickness with halo structural parameters, spin and concentration, which are the most important factors after mass and redshift according to the SHAP scores. Here we fix the halo mass to the Milky-Way scale ($M_{\rm vir}=10^{11.7-12.3}M_{\odot}$) to better focus on the trend with the structural parameters. Disk thickness increases with spin for both hydro halos and DMO halos, while a dependence on concentration appears only in the hydro case. Interestingly, once disk compactness is fixed, as for the lower panels of Fig. 10, these correlations largely disappear. This suggests that spin and concentration affect thickness indirectly – through their impact on disk size and compactness – rather than by directly regulating the vertical velocity dispersion.
Finally, we examine disk thickness versus halo assembly–history parameters in Fig. 11. At fixed halo mass, lower formation redshift $z_{1/2}$ and lower normalized accretion rate $\dot{M}_{\rm norm}$ – corresponding to more recent or more rapid mass growth – produce systematically thicker disks at all redshifts. This is a natural consequence of dynamical heating from mergers and accretion. Taken together, these results indicate that active mass assembly is the primary driver of disk thickness, whereas the apparent correlations with halo spin and concentration arise indirectly through their shared impact on disk compactness. Although a full explanation of how halo structure governs disk compactness and why disk size and height correlate so tightly lie beyond the scope of this work, these results provide a clear clue as to why halo properties predict disk thickness better than disk size. The additional predictive power comes from mass-assembly parameters, which exert a stronger influence on disk thickness than on disk compactness.
4.5 Do inner halo properties improve disk–halo modeling?
As shown earlier in Section 3, inner-halo parameters emerge as highly important predictors of disk structure—but only when measured from the hydrodynamical simulation. To clarify this difference, Fig. 12 compares the predictive power of the global spin $\lambda$ and the inner-halo spin $\lambda_{\rm inner}$ for both disk compactness and disk thickness.
@@ 4.5 Do inner halo properties improve disk–halo modeling?
4.6 Model applicability to subsamples
One may ask whether the regression relations remain robust when the subgrid physics prescriptions are varied. In principle, this could be tested by applying the models to a different simulation suite with alternative subgrid implementations. While such a comparison is beyond the scope of the current work, we can perform a related internal test within the TNG sample by examining whether the models hold for subsamples selected by specific star formation rate (sSFR). A primary effect of modifying subgrid prescriptions is to alter the global star-formation activity of galaxies. Therefore, if the model performs well for subsamples split by sSFR, this would suggest that it is relatively insensitive to such recipe variations.
@@ 4.6 Model applicability to subsamples
5 Conclusion

### v2

@@ III.2.1 Disk size
Feature importance: The importance rankings differ between the hydro and DMO cases. In the hydro case, the Einasto shape index $\alpha$ is the dominant factor for disk size, while it does not even appear in the top 10 in the DMO case. Some features are consistently influential for both hydro and DMO halos, including concentration $c$, spin $\lambda$, virial mass $M_{\rm vir}$, redshift $z$, and inner-halo spin $\lambda_{\rm inner}$. The most important halo properties align with those used in previous disk-size models (e.g. Mo et al., 1998; Jiang et al., 2019). Inner-halo parameters frequently emerge as important features when using the hydro halos, whereas global halo parameters tend to be more important when using the DMO halos. This difference arises mainly because the disk interacts via gravity with the surrounding DM that constitutes the inner halo, altering its structure, as will be further illustrated in Section IV.5. Halo shape parameters are always among the top 10 features. As discussed in Jiang et al. (2025) and Liang et al. (2025), this is likely because rounder shapes are indicative of more relaxed halos in which extended disks can more easily develop. Environment and assembly-history parameters (e.g., $n_{1}$, PC1, PC2, $\dot{M}_{\rm norm}$) have limited impact.
@@
III.2.2 Disk scale height
@@
III.2.3 Disk mass fraction
@@
Figure 6: Comparison of models relating disk size to halo properties. Disk half-mass radii $R_{\rm 1/2,Disk}$ measured from the TNG50 hydro simulation are compared with predictions from several models.The first panel shows our SR-based model, written for simplicity as $R_{\rm 1/2,Disk}=f(X)R_{\rm vir}$, where $f(X)$ denotes the SR formula as a function of halo parameters $X$. The cyan and red curves correspond to the SR relations calibrated on hydro halos and DMO halos, respectively, with the DMO version given by eq. (21) and the hydro version by $f(X)=\lambda_{\rm inner}+(\lambda+c^{-0.125})(\log\alpha+3.010)-3.210$. The second panel shows the relation of Kravtsov (2013), $R_{1/2}=0.02R_{\rm vir}$.The third panel shows a simplified version of the spin-based model of Mo et al. (1998), $R_{1/2}=0.5\lambda R_{\rm vir}$.The fourth panel shows the concentration-based empirical model of Jiang et al. (2019), $R_{1/2}=0.02(1+z)^{-0.2}(c/10)^{-0.7}R_{\rm vir}$. Median relations using hydro (DMO) halos are shown as solid cyan (dashed red) lines, with shaded regions indicating the 16–84th percentiles. Spearman correlation coefficients are listed in each panel, and the diagonal line marks the one-to-one relation. This analysis is limited to the galaxies with significant disks ($f_{\rm Disk}>0.3$).  Our SR models provide the highest correlation and lowest scatter, outperforming all previous prescriptions.
@@ IV Discussion
IV.1 New size predictors based on symbolic regression
Simple analytic prescriptions have long been used to link galaxy size to halo properties. Kravtsov (2013) derived a linear relation $R_{1/2}=0.02\,R_{\rm vir}$ from abundance matching, applicable to both disks and spheroids. Mo et al. (1998) modeled thin exponential disks that inherit the halo’s specific angular momentum, yielding a size–spin scaling. Following Somerville et al. (2018), we adopt a much simplified form of the original Mo et al. model, $R_{1/2}=0.5\,\lambda\,R_{\rm vir}$. In contrast, Jiang et al. (2019) found marginal correlation between halo spin and galaxy size in cosmological zoom-in simulations and proposed a concentration-based scaling, $R_{1/2}=0.02(1+z)^{-0.2}(c/10)^{-0.7}R_{\rm vir}$.
@@ IV.1 New size predictors based on symbolic regression
IV.2 Understanding the relation between disk size and Einasto shape index
@@
In Section III, the SHAP analysis shows that the density–profile shape index, $\alpha$, is a key predictor of disk compactness, $R_{1/2,\rm Disk}/R_{\rm vir}$ – but only when using halo properties measured from the hydrodynamic simulation. In the DMO case, $\alpha$ carries far less predictive power. This contrast suggests that the apparent link between disk size and $\alpha$ arises from baryonic effects, rather than from the primordial structure of DMO halos. This is further explored in Fig. 7, which shows a clear positive correlation between $R_{1/2,\rm Disk}/R_{\rm vir}$ and $\alpha$ in the hydro halos, but almost no correlation or weakly negative correlation in the DMO counterparts.
@@
which quantifies the halo’s structural response to baryonic processes. Here, $\alpha_{\rm Hydro}$ is the Einasto index measured in the hydro halo, and $\alpha_{\rm DMO}$ is that of the matched DMO halo. By construction, $\tilde{\alpha}<0$ corresponds to halo contraction (Blumenthal et al., 1986; Gnedin et al., 2004), whereas $\tilde{\alpha}>0$ indicates expansion. We also examine the stellar-to-halo mass ratio, $M_{\star}/M_{\rm vir}$, as a proxy for the strength of baryonic impact (Di Cintio et al., 2014; Freundlich et al., 2020). As shown in the right panels of Fig. 7, most halos have $\tilde{\alpha}<0$, demonstrating that contraction is the dominant response in TNG50. Moreover, the mean value of $\tilde{\alpha}$ decreases with increasing $M_{\star}/M_{\rm vir}$, indicating that halos in more massive galaxies tend to contract more strongly. At fixed stellar-to-halo mass ratio – for example, around $M_{\star}/M_{\rm vir}\simeq 0.03$, where systems are comparably disk-dominated – we find a clear trend that disk compactness governs the degree of halo contraction. The most extended disks (the yellow points in Fig. 7) can even drive halo expansion rather than contraction.
IV.3 Redshift trend of disk compactness
@@
To understand these redshift trends, we first focus on the low-mass regime and ask what drives the strong evolution with redshift. A simple and intuitive picture emerges when one approximates the disk size as the radius at which gas first crosses the star-formation density threshold, and considers how the virial radius evolves at fixed halo mass. In the TNG model, star formation begins once the hydrogen number density exceeds $0.13\mathrm{cm}^{-3}$. Using a toy model with an NFW halo and a self-similar gas distribution containing the cosmic baryon fraction, we find that this star-forming radius remains roughly constant across redshift, primarily because typical halo concentrations decrease toward high $z$. By contrast, the virial radius, $R_{\rm vir}\propto(M_{\rm vir}/\rho_{\rm crit})^{1/3}$, becomes much smaller at early times due to the rapid rise of the critical density. As a result, even if the physical disk size evolves only weakly, the ratio $R_{1/2,\rm Disk}/R_{\rm vir}$ increases steeply toward high redshift. This picture is most applicable to low-mass halos, which predominantly grow via cold-mode accretion (Dekel et al., 2009), where cooling is efficient and disk size closely tracks the extent of star-forming gas.
For more massive systems ($M_{\rm vir}\gtrsim 2\times 10^{11}M_{\odot}$), the processes that shape disk size differ markedly with redshift. At high $z$, these galaxies experience rapid mass assembly and frequent gas-rich mergers, which trigger centrally concentrated starbursts and produce compact stellar components (Zolotov et al., 2015; Lapiner et al., 2023). Although often described as star-forming nuggets with spheroidal morphologies, these systems can still retain substantial angular momentum inherited from merger orbits or inflowing filaments. Because our morphological decomposition identifies disks via the angular-momentum distribution of stellar particles, rapidly rotating nuggets are naturally classified as disks. This is not specific to our method – most kinematics-based decomposition schemes behave similarly (e.g., Du et al., 2019; Zana et al., 2022) – reflecting the broader fact that high-$z$ disks and fast rotators are not cleanly separable. Consequently, the rotating stellar components in massive high-redshift halos tend to be compact.
At low redshift ($z\lesssim 1$), by contrast, galaxy growth is more quiescent and proceeds inside-out (Genel et al., 2018). Halos of a given mass are also more dynamically settled, providing favorable conditions for the formation of extended disks (Jiang et al., 2025). This naturally explains the monotonic increase of disk size with halo mass at low redshift.
IV.4 Drivers of disk thickness: mass, size, and halo growth history
As shown in Fig. 4 and discussed in Section III.2.2, halo properties exhibit unexpectedly strong predictive power for disk thickness, stronger, in fact, than for disk compactness. Here we explore clues to the possible physical origin of this behavior.
@@ IV.4 Drivers of disk thickness: mass, size, and halo growth history
The disk flattening ratio, $Z_{\rm 1/2,Disk}/R_{\rm 1/2,Disk}$, reaches a minimum of $\sim 0.1$ near $M_{\rm vir}\sim 10^{12}M_{\odot}$ and rises at both lower and higher masses. At any fixed mass, disks are systematically thicker at earlier cosmic times. This potentially reflects enhanced turbulence and elevated vertical velocity dispersion in high-redshift galaxies, consistent with vertical Jeans equilibrium, where $Z_{1/2}\propto\sigma_{z}^{2}/(\pi G\Sigma)$.
@@
A second key finding is the tight correlation between disk height and disk size. As shown in the middle and right panels of Fig. 9, disk height increases almost linearly with disk size at fixed halo mass. This implies that many halo properties might influence disk thickness primarily through their regulation of disk size—so the factors that set disk extent also tend to set disk thickness.
@@
We then examine in Fig. 10 the correlations of thickness with halo structural parameters, spin and concentration, which are the most important factors after mass and redshift according to the SHAP scores. Here we fix the halo mass to the Milky-Way scale ($M_{\rm vir}=10^{11.7-12.3}M_{\odot}$) to better focus on the trend with the structural parameters. Disk thickness increases with spin for both hydro halos and DMO halos, while a dependence on concentration appears only in the hydro case. Interestingly, once disk compactness is fixed, as for the lower panels of Fig. 10, these correlations largely disappear. This may suggests that spin and concentration affect thickness indirectly – through their impact on disk size and compactness – rather than by directly regulating the vertical velocity dispersion.
Finally, we examine disk thickness versus halo assembly–history parameters in Fig. 11. At fixed halo mass, lower formation redshift $z_{1/2}$ and lower normalized accretion rate $\dot{M}_{\rm norm}$ – corresponding to more recent or more rapid mass growth – produce systematically thicker disks at all redshifts. This is a natural consequence of dynamical heating from mergers and accretion. Taken together, these results might indicate that active mass assembly is the primary driver of disk thickness, whereas the apparent correlations with halo spin and concentration arise indirectly through their shared impact on disk compactness. Although a full explanation of how halo structure governs disk compactness and why disk size and height correlate so tightly lie beyond the scope of this work, these results provide a clear clue as to why halo properties predict disk thickness better than disk size. The additional predictive power possibly comes from mass-assembly parameters, which exert a stronger influence on disk thickness than on disk compactness.
IV.5 Do inner halo properties improve disk–halo modeling?
As shown earlier in Section III, inner-halo parameters emerge as highly important predictors of disk structure—but only when measured from the hydrodynamical simulation. To clarify this difference, Fig. 12 compares the predictive power of the global spin $\lambda$ and the inner-halo spin $\lambda_{\rm inner}$ for both disk compactness and disk thickness.
@@ IV.5 Do inner halo properties improve disk–halo modeling?
IV.6 Generalizability beyond TNG physics
Our results are based on the TNG simulations. Therefore, the feature importance identified by the RF and SHAP analyses, the empirical relations learned by SR, and the correlations are specific to TNG physics rather than universally applicable. For example, halo spin parameters emerge as key predictors of disk structure in our analysis. Similarly strong correlations with spin are also reported in Paper I and in Liao et al. (2019) using the Auriga simulation (Grand et al., 2017). However, other simulations may show weaker spin dependence. Rohr et al. (2022) based on FIREbox (Feldmann et al., 2023), Jiang et al. (2019) based on NIHAO (Wang et al., 2015) and VELA (Ceverino et al., 2014; Zolotov et al., 2015) find that halo spin is not strongly linked to galaxy size. Yang et al. (2023) demonstrated that correlations between galaxy compactness and halo spin are significantly stronger in TNG than in EAGLE. We speculate that these differences arise from more efficient baryonic cycles, and will follow up in a future study.
@@ IV.6 Generalizability beyond TNG physics
Therefore, applying the same analytical framework to those simulations may not highlight the spin parameter as an important factor, and the SR relations may not perform well using halo properties from those simulations.
To test whether the regression relations remain robust when the subgrid physics prescriptions are varied requires applying the models to a different simulation suite. While such a comparison is beyond the scope of the current work, we can perform a related internal test within the TNG sample by examining whether the models hold for subsamples selected by specific star formation rate (sSFR). A primary effect of modifying subgrid prescriptions is to alter the global star-formation activity of galaxies. Therefore, if the model performs well for subsamples split by sSFR, this would suggest that it is relatively insensitive to such recipe variations.
@@
IV.7 Resolution effects on disk structures
To test whether disk structures (e.g. size and height) are well resolved, we examine the distributions of disk size and disk height and compare them with the collisionless softening length (0.288 ckpc). For disk size, the 16th percentile is above the softening length for all samples at all redshifts. For disk height, the 16th percentile for total samples at $z=0$ and for MW-mass systems at $z=2$ and $z=4$ is comparable to the softening length, indicating marginally sufficient resolution for some cases. We experimented with restricting our analysis to well-resolved subsamples, and find similarly high prediction accuracy for both our RF and SR models, suggesting that the ML models might capture underlying physical correlations rather than being sensitive to numerical resolution effects. However, applying our pipeline to higher resolution simulations might still be needed since some cases are marginally resolved.
IV.8 Sample completeness at low masses
Following Paper I, we require a minimum of 1000 stellar and 1000 DM particles. The threshold on the number of stellar particles ensures robust morphological decomposition but is the more restrictive condition of the two, since galaxies with 1,000 stellar particles typically reside in halos well above 1000 DM particles. For this reason, our sample and the measurements for halo quantities becomes incomplete at the low-mass end. To qualitatively describe this bias: galaxies that are strongly DM-dominated are missing from our analysis at the low-mass end.
IV.9 The selection of best-fitting functions
The SR requires a metric to select functions at each iteration, with MSE-like losses (e.g. Eq. 11, Eq. II.4) being the simplest choice. However, since such losses typically decrease with increasing complexity, model selection must balance accuracy and simplicity, often via heuristic penalties or Pareto fronts. These approaches may under-penalize complexity. Information criteria such as Akaike (AIC), Bayesian (BIC), and Deviance (DIC) provide an alternative by penalizing parameter count, but they do not capture functional complexity. A more principled approach is the minimum description length (MDL) principle (e.g. Grünwald and Roos, 2019), which jointly quantifies accuracy and complexity from an information-theoretic perspective and has been applied in astrophysical SR studies (e.g. Bartlett et al., 2024; Desmond, 2025). Here we adopt MSE-like losses for simplicity and good empirical performance, while encouraging future work to explore metrics that better penalize functional complexity.
V Conclusion